# Rigorous dense graph limit of a model for

### v1

@@
Abstract. We rigorously derive the dense graph limit of a discrete model describing the formation of biological transportation networks. The discrete model, defined on undirected graphs with pressure-driven flows, incorporates a convex energy functional combining pumping and metabolic costs. It is constrained by a Kirchhoff law reflecting the local mass conservation. We first rescale and reformulate the discrete energy functional as an integral ‘semi-discrete’ functional, where the Kirchhoff law transforms into a nonlocal elliptic integral equation. Assuming that the sequence of graphs is uniformly connected and that the limiting graphon is $0-1$ valued, we prove two results: (1) rigorous $\Gamma$-convergence of the sequence of the semi-discrete functionals to a continuum limit as the number of graph nodes and edges tends to infinity; (2) convergence of global minimizers of the discrete functionals to a global minimizer of the limiting continuum functional. Our results provide a rigorous mathematical foundation for the continuum description of biological transport structures emerging from discrete networks.
@@ 1. Introduction
In this paper we derive the rigorous dense graph limit of the discrete network formation model introduced and studied in [14, 15]. The model is posed on a given connected graph $\mathbb{G}=(\mathbb{V},\mathbb{E})$, consisting of a set of vertices (nodes) $\mathbb{V}$ and a set of unoriented edges (vessels) $\mathbb{E}\subset\mathbb{V}\times\mathbb{V}$. Any two vertices $i$, $j\in\mathbb{V}$ are linked by at most one edge $(i,j)\in\mathbb{E}$, and we also assume that there are no loops, i.e., $(i,i)\notin\mathbb{E}$ for all $i\in\mathbb{V}$. Since the graph is undirected we refer by $(i,j)$ and $(j,i)$ to the same edge. The lengths $L_{ij}=L_{ji}>0$ of the edges $(i,j)\in\mathbb{E}$ are given and fixed. Moreover, we denote by $W_{ij}=W_{ji}$ its adjacency matrix, i.e., $W_{ij}=1$ if $(i,j)\in\mathbb{E}$ and $W_{ij}=0$ otherwise.
@@
$-\sum_{j\in\mathcal{\mathbb{V}}}W_{ij}C_{ij}\frac{P_{j}-P_{i}}{L_{ij}}=S_{i}\qquad\mbox{for all }i\in\mathbb{V}.$ (1.3)
@@
$(P_{j}-P_{i})Q_{ij}=\frac{Q_{ij}^{2}}{C_{ij}}L_{ij}.$
The metabolic cost of maintaining the edge $(i,j)\in\mathbb{E}$ is assumed to be proportional to its length $L_{ij}$ and a power of its conductivity $C_{ij}^{\gamma}$, with the metabolic exponent $\gamma>0$. In this paper we shall focus on the case $\gamma>1$. Consequently, the discrete energy cost functional is given by
$E[C]:=\frac{1}{2}\sum_{i\in\mathbb{V}}\sum_{j\in\mathbb{V}}\left(C_{ij}\frac{(P_{j}-P_{i})^{2}}{L_{ij}^{2}}+\frac{\nu}{\gamma}C_{ij}^{\gamma}\right)W_{ij}L_{ij}$ (1.4)
where the pressures $P_{ij}=P_{ij}[C]$ are determined as a solution to Kirchhoff’s law (1.3), and $\nu>0$ is the so-called metabolic coefficient. If Kirchhoff’s law does not admit a solution with the given matrix $C=(C_{ij})_{(i,j)\in\mathbb{V}\times\mathbb{V}}$, we set $E[C]:=+\infty$.
@@
[…]
@@
In this paper we derive the graphon limit of (1.3)-(1.4), where the limiting integral energy functional is defined on the set of nonnegative scalar functions. These scalar functions represent network conductivities between nodes of the limiting graphon and are defined on the square $[0,1]^{2}$, where the interval $[0,1]$ represents node labels. The energy functional is constrained by an integral Poisson-type equation for the pressure. The gradient of the pressure, which appears in the “classical” Poisson equation, is replaced by finite pressure difference between the corresponding nodes. The scalar conductivity, multiplied by the graphon function, acts as the integral kernel. Let us stress that our derivation of the limiting functional as a $\Gamma$-limit of (1.3)-(1.4) is rigorous, i.e., not only formal, as in [9] and [11].
@@ 2. Notation and preliminaries
$\mathcal{I}_{i}^{N}:=\left[\frac{i-1}{N},\frac{i}{N}\right]\qquad\mbox{for }i\in[N]$
where here and in the sequel we use the notation $[N]:=\{1,2,\dots,N\}$. Moreover, for $N\in\mathbb{N}$ we denote
@@
Moreover, for $q\in[1,+\infty]$ and $r>0$ we introduce the convex set
@@
$\mathbb{B}^{N}_{r}:=\left\{B\in\mathbb{R}^{N\times N}_{\mathrm{sym}}\,\middle|\,\begin{array}[]{l}B_{ij}\geq r\quad\text{for all }(i,j)\in[N]^{2}\text{ such that }W^{N}_{ij}=1,\\ B_{ij}=0\quad\text{for all }(i,j)\in[N]^{2}\text{ such that }W^{N}_{ij}=0\end{array}\right\}.$ (2.4)
@@
$\mathfrak{d}p(x,y):=p(x)-p(y)\qquad\mbox{for }(x,y)\in\mathcal{I}\times\mathcal{I}.$ (2.5)
@@
where $\chi^{N}_{i}=\chi_{I^{N}_{i}}$ is the characteristic function of the interval $I^{N}_{i}$. By a slight abuse of notation, we also use the symbol $\mathcal{Q}^{N}$ for the operator $\mathcal{Q}^{N}:\mathbb{R}^{N\times N}\to L^{\infty}(\mathcal{I}\times\mathcal{I})$, mapping matrices $B\in\mathbb{R}^{N\times N}$ onto piecewise constant functions on $\mathcal{I}\times\mathcal{I}$,
@@ 3. Main assumptions
$w^{N}\to w\mbox{ in the norm topology of }L^{1}(\mathcal{I}\times\mathcal{I})\mbox{ as }N\to+\infty.$ (3.3)
@@
Assumption (3.3) restricts the validity of our results to $0-1$ valued graphons, i.e., $\mbox{range}(w)\subset\{0,1\}$. In this case one can show [19, Proposition 8.24] that the usual convergence of $w^{N}$ to $w$ in the cut norm indeed implies convergence in the $L^{1}$ norm topology. Of course, due to the uniform boundedness of $w^{N}$ in $L^{\infty}(\mathcal{I}\times\mathcal{I})$, the convergence (3.3) holds also in the norm topology of $L^{q}(\mathcal{I}\times\mathcal{I})$ for any $q<+\infty$. Let us explicitly note that (3.3) excludes important classes of graphs. For instance, for the generic random Erdös-Rényi graph model [4], where pair of nodes are connected with probability $p\in(0,1)$, we have $w\equiv p$ on $\mathcal{I}\times\mathcal{I}$. Obviously, it is then not possible that $w^{N}\to w$ strongly in $L^{1}$, since the functions $w^{N}$ are piecewise constant taking only values 00 and $1$. We shall generalize our theory to include random graph models in a future work.
To be able to pass to the continuum limit in (1.3)-(1.4), we also need to assume that the edge lengths $L^{N}=(L^{N}_{ij})_{i,j=1}^{N}$ converge in an appropriate sense. In particular, for $\ell^{N}:=\mathcal{Q}^{N}[L^{N}]$ we impose the assumption that
$\left\|\ell^{N}\right\|_{L^{\infty}(\mathcal{I}\times\mathcal{I})}=\sup_{i,j\in[N]}L^{N}_{ij}\leq 1,$ (3.4)
@@
$\ell^{N}\to\ell\mbox{ in the norm topology of }L^{1}(\mathcal{I}\times\mathcal{I})\mbox{ as }N\to+\infty.$ (3.5)
Assumption (3.4) is a natural assumption meaning that the sequence of graphs $\mathbb{G}^{N}$ is located within a compact set in the “physical space”. The particular bound $L_{ij}\leq 1$ can be achieved through rescaling and is therefore assumed without loss of generality. A trivial way to ensure validity of assumption (3.5) is to fix $\ell\in L^{\infty}(\mathcal{I}\times\mathcal{I})$ and set $\ell^{N}:=\mathcal{Z}^{N}[\ell]$. The lengths $L^{N}_{ij}$ are then the corresponding values of $\ell^{N}$, i.e., such that $\ell^{N}:=\mathcal{Q}^{N}[L^{N}]$ holds. Again, due to the uniform boundedness of $\ell^{N}$ in $L^{\infty}(\mathcal{I}\times\mathcal{I})$, the limit in (3.5) holds also with respect to the norm topology of $L^{q}(\mathcal{I}\times\mathcal{I})$ for any $q<+\infty$.
Assumptions (3.2)–(3.5) are sufficient for proving our first main result, which establishes the $\Gamma$-convergence of the sequence of (properly rescaled and reformulated) energy functionals. For our second main result, which states that the sequence of global minimizers of the discrete functionals (1.3)-(1.4) converges to the global minimizer of the limiting functional, we need to adopt one more assumption. Namely, we need to assume that the sequence of reciprocal values $\left(\ell^{N}\right)^{-1}$ is uniformly bounded in $L^{\frac{2(\gamma+1)}{\gamma}}(\mathcal{I}\times\mathcal{I})$,
$\sup_{N\in\mathbb{N}}\left\|\left(\ell^{N}\right)^{-1}\right\|_{L^{\frac{2(\gamma+1)}{\gamma}}(\mathcal{I}\times\mathcal{I})}<+\infty.$ (3.6)
Admittedly, this is a rather strong assumption and constitutes a key limitation of our approach. To gain an intuition about the satisfiability of (3.6), let us consider the case when the vertices in $\mathbb{V}^{N}$ are generated as random points in a convex compact set $K\subset\mathbb{R}^{d}$. I.e., we identify $\mathbb{V}^{N}$ with a cloud of $N\in\mathbb{N}$ points drawn independently from the uniform distribution on $K$. The edge lengths $L_{ij}$ are then simply the Euclidean distances between the points (vertices) $i$ and $j$. The law of large numbers [2] gives then
$\lim_{N\to\infty}\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j\neq i}^{N}L_{ij}^{-q}=\int_{0}^{+\infty}s^{-q}f(s)\,\mathrm{d}s,$ (3.7)
where $f=f(s)$ is the probability density for the Euclidean distances of the point cloud.
For the case when $K$ is a hypercube in $\mathbb{R}^{d}$, the probability density $f=f(s)$ is known explicitly, see, e.g., [16, 17] and references therein. For $d=2$ and $K=[0,1/\sqrt{2}]^{2}$, we have $L_{ij}\leq 1$ and
@@
complemented by a more complicated formula for values of $s\in\left(\frac{1}{\sqrt{2}},1\right]$. However, $f(s)$ is decreasing for $s>1/\sqrt{2}$, see [17], so that we have the bound $f(s)<f\left(1/\sqrt{2}\right)=2\sqrt{2}(\pi-3)$ for $s>1/\sqrt{2}$. Of course, $f(s)=0$ for all $s>1$. Then, in (3.7) we have
@@
Obviously, the right-hand side is finite if and only if the first integral is finite, which holds if and only if $q<2$. Unfortunately, this makes assumption (3.6) unsatisfiable since $\frac{2(\gamma+1)}{\gamma}>2$ for all $\gamma>0$.
In general, the probability density for the $d$-dimensional cube behaves like $s^{d-1}$ as $s\to 0+$, [17]. Consequently, (3.7) is finite if and only if $q<d$. For $d=3$ the assumption (3.6) means $\frac{2(\gamma+1)}{\gamma}<3$, which is equivalent to $\gamma>2$. Finally, for $d=4$ (or, eventually, higher) we can impose (3.6) for any $\gamma>1$.
@@ 4. Reformulation of the model and main results
For the purpose of deriving the graphon limit of the model (1.3)-(1.4), we first introduce a regularization by adding a fixed constant $r>0$ to all conductivities $C_{ij}$. This is necessary to ensure uniform ellipticity of the Kirchhoff law (1.3) and was also introduced in the previous works [9, 10] and [11] where continuum limits were derived. It is then convenient to introduce the new set of variables $B_{ij}:=\frac{C_{ij}+r}{L_{ij}}$. Moreover, the energy functional $E^{N}$ has to be scaled by $1/N^{2}$, which can be interpreted as the reciprocal of the square of the “mean” edge length. This is due to the fact that we are embedding the graph $\mathbb{G}^{N}$, which is inherently a one-dimensional structure, into the two-dimensional interval $\mathcal{I}\times\mathcal{I}$. The same scaling was adopted in [9, 10, 11] in order to obtain a meaningful continuum limit. We therefore introduce the rescaled energy functional
$F^{N}[B]:=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\left(B_{ij}(P_{j}-P_{i})^{2}+\frac{\nu}{\gamma}B_{ij}^{\gamma}\left(L_{ij}^{N}\right)^{\gamma+1}\right)W^{N}_{ij},$ (4.1)
@@
Clearly, we have $F^{N}[B]=E^{N}[C+r]/N^{2}$, with $(C+r)_{ij}:=C_{ij}+r$. We fix $r>0$ and, recalling the assumption $L_{ij}\leq 1$, pose (4.1)–(4.2) on the convex set $\mathbb{B}^{N}_{r}$ defined in (2.4). We shall also use the “weak formulation” of (4.2), which is obtained by its multiplication by $\Phi_{i}$ for a “test vector” $\Phi\in\mathbb{R}_{0}^{N}$ and summation over $i\in[N]$,
@@
where we use the concise notation $\,\mathrm{d}w(x,y)$ for $w(x,y)\,\mathrm{d}x\,\mathrm{d}y$ and $w\in L^{\infty}(\mathcal{I}\times\mathcal{I})$ is given in (3.3). The functional $\mathcal{F}$ is defined on the set $L^{\infty}_{r}(\mathcal{I}\times\mathcal{I})$ introduced in (2.1). The scalar pressure $p=p[b]\in L^{2}_{0}(I)$ is obtained as a solution of the nonlocal diffusion (Poisson-type) equation
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)(\mathfrak{d}p(x,y))(\mathfrak{d}\varphi(x,y))\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma(x)\varphi(x)\,\mathrm{d}x\quad\mbox{for all }\varphi\in L^{2}_{0}(\mathcal{I}),$ (4.5)
where we used the notation (2.5). Let us note that the functional (4.4)–(4.5) is invariant with respect to any measure preserving change of variable on the interval $\mathcal{I}$. Indeed, if $\psi:\mathcal{I}\to\mathcal{I}$ is any such measure preserving bijection, then the value of the energy remains the same for the transformed functions $\tilde{b}(x,y):=b(\psi(x),\psi(y))$, ~$\tilde{\ell}(x,y):=\ell(\psi(x),\psi(y))$, $\tilde{w}(x,y):=w(\psi(x),\psi(y))$ and $\tilde{p}(x):=p(\psi(x))$, $\tilde{\sigma}(x):=\sigma(\psi(x))$. This invariance is important since $w$ and $\tilde{w}$ represent the same graphon [3, 19]. An analogous invariance obviously also holds for the discrete model (4.1)–(4.2) with respect to permutations of the set of indices $[N]$.
@@
$\mathcal{F}^{N}[b]:=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\left(\mathfrak{d}p^{N}(x,y)\right)^{2}+\frac{\nu}{\gamma}b(x,y)^{\gamma}\ell^{N}(x,y)^{\gamma+1}\,\mathrm{d}w^{N}(x,y),$ (4.6)
with $\ell^{N}=\mathcal{Q}^{N}[L^{N}]$, $w^{N}=\mathcal{Q}^{N}[W^{N}]$, and $p^{N}\in L^{2}_{0,N}(\mathcal{I})$ is the unique piecewise constant solution of the approximate Poisson equation
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)(\mathfrak{d}p^{N}(x,y))(\mathfrak{d}\varphi^{N}(x,y))\,\mathrm{d}w^{N}(x,y)=\int_{\mathcal{I}}\sigma(x)\varphi^{N}(x)\,\mathrm{d}x\quad\mbox{for all }\varphi^{N}\in L^{2}_{0,N}(\mathcal{I}).$ (4.7)
@@
Moreover, the sequence of functionals $(\mathcal{F}^{N})_{N\in\mathbb{N}}$, given by (4.6)–(4.7), $\Gamma$-converges in $L^{\omega}(\mathcal{I}\times\mathcal{I})$, $\omega=\max\{2,\gamma\}$, to the functional $\mathcal{F}$ given by (4.4)–(4.5), in the following sense:

### v2

@@
Abstract. We rigorously derive the dense graph limit of a discrete model describing the formation of biological transportation networks. The discrete model, defined on undirected graphs with pressure-driven flows, incorporates a convex energy functional combining pumping and metabolic costs. It is constrained by a Kirchhoff law reflecting the local mass conservation. We first rescale and reformulate the discrete energy functional as an integral ‘semi-discrete’ functional, where the Kirchhoff law transforms into a nonlocal elliptic integral equation. Assuming that the sequence of graphs is uniformly connected and that the limiting graphon is 010-1 valued, we prove two results: (1) rigorous $\Gamma$-convergence of the sequence of the semi-discrete functionals to a continuum limit as the number of graph nodes and edges tends to infinity; (2) convergence of global minimizers of the discrete functionals to a global minimizer of the limiting continuum functional. Our results provide a rigorous mathematical foundation for the continuum description of biological transport structures emerging from discrete networks.
@@ 1. Introduction
In this paper we derive the rigorous dense graph limit of the discrete network formation model introduced and studied in [14, 15]. The model is posed on a given connected graph $\mathbb{G}=(\mathbb{V},\mathbb{E})$, consisting of a set of vertices (nodes) $\mathbb{V}$ and a set of unoriented edges (vessels) $\mathbb{E}\subset\mathbb{V}\times\mathbb{V}$. Any two vertices $i$, $j\in\mathbb{V}$ are linked by at most one edge $(i,j)\in\mathbb{E}$, and we also assume that there are no loops, i.e., $(i,i)\notin\mathbb{E}$ for all $i\in\mathbb{V}$. Since the graph is undirected we refer by $(i,j)$ and $(j,i)$ to the same edge. The lengths $\mathrm{L}_{ij}=\mathrm{L}_{ji}>0$ of the edges $(i,j)\in\mathbb{E}$ are given and fixed. Moreover, we denote by $W_{ij}=W_{ji}$ its adjacency matrix, i.e., $W_{ij}=1$ if $(i,j)\in\mathbb{E}$ and $W_{ij}=0$ otherwise.
@@
$-\sum_{j\in\mathcal{\mathbb{V}}}W_{ij}C_{ij}\frac{P_{j}-P_{i}}{\mathrm{L}_{ij}}=S_{i}\qquad\mbox{for all }i\in\mathbb{V}.$ (1.3)
@@
$(P_{j}-P_{i})Q_{ij}=\frac{Q_{ij}^{2}}{C_{ij}}\mathrm{L}_{ij}.$
The metabolic cost of maintaining the edge $(i,j)\in\mathbb{E}$ is assumed to be proportional to its length $\mathrm{L}_{ij}$ and a power of its conductivity $C_{ij}^{\gamma}$, with the metabolic exponent $\gamma>0$. In this paper we shall focus on the case $\gamma>1$. Consequently, the discrete energy cost functional is given by
$E[C]:=\frac{1}{2}\sum_{i\in\mathbb{V}}\sum_{j\in\mathbb{V}}\left(C_{ij}\frac{(P_{j}-P_{i})^{2}}{\mathrm{L}_{ij}^{2}}+\frac{\nu}{\gamma}C_{ij}^{\gamma}\right)W_{ij}\mathrm{L}_{ij}$ (1.4)
where the pressures $P_{ij}=P_{ij}[C]$ are determined as a solution to Kirchhoff’s law (1.3), and $\nu>0$ is the so-called metabolic coefficient. If Kirchhoff’s law does not admit a solution with the given matrix $C=(C_{ij})_{(i,j)\in\mathbb{V}\times\mathbb{V}}$, we set $E[C]:=\infty$.
@@
A graphon is a symmetric measurable function $w:[0,1]^{2}\to[0,1]$. In the context of our model (1.3)-(1.4) we shall interpret $w=w(x,y)$ as the limit of the sequence of graphs $\mathbb{G}^{N}=(\mathbb{V}^{N},\mathbb{E}^{N})$ as $N\to\infty$, where $|\mathbb{V}^{N}|=N$. The goal of this paper is then to establish the rigorous limit passage of the energy functional $E=E[C]$, given by (1.3)-(1.4), as $N\to\infty$. In the first step we shall rescale and reformulate the energy functional in terms of an integral functional acting on the space of piecewise constant functions on $0,1]^{2}$. We then establish its limit in the sense of $\Gamma$-convergence [20] to an integral functional, where the Kirchhoff law (1.3) becomes an elliptic integral equation in the limit. The $\limsup$-condition’ of $\Gamma$-convergence follows from strong continuity of the limiting functional with respect to the appropriate topology. The $\liminf$-condition’ is then a consequence of its weak lower semicontinuity, which follows from convexity (for values of the metabolic exponent $\gamma\geq 1$). Finally, using the $\Gamma$-convergence, we shall conclude that global minimizers of the discrete energy functional (1.3)-(1.4) converge, in appropriate sense, to global minimizers of the limiting (continuum) functional.
@@
In this paper we derive the graphon limit of (1.3)-(1.4), where the limiting integral energy functional is defined on the set of nonnegative scalar functions. These scalar functions represent network conductivities between nodes of the limiting graphon and are defined on the square $0,1]^{2}$, where the interval $[0,1]$ represents node labels. The energy functional is constrained by an integral Poisson-type equation for the pressure. The gradient of the pressure, which appears in the “classical” Poisson equation, is replaced by finite pressure difference between the corresponding nodes. The scalar conductivity, multiplied by the graphon function, acts as the integral kernel. Let us stress that our derivation of the limiting functional as a $\Gamma$-limit of (1.3)-(1.4) is rigorous, i.e., not only formal, as in [9] and [11].
@@ 2. Notation and preliminaries
[…]
where here and in the sequel we use the notation $N]:=\{1,2,\dots,N\}$. Moreover, for $N\in\mathbb{N}$ we denote
@@
Moreover, for $q\in[1,\infty]$ and $r>0$ we introduce the convex set
@@
[…]
@@
$\overline{\nabla}p(x,y):=p(y)-p(x)\qquad\mbox{for }(x,y)\in\mathcal{I}\times\mathcal{I}.$ (2.5)
@@
where $\chi^{N}_{i}=\chi_{\mathcal{I}^{N}_{i}}$ is the characteristic function of the interval $\mathcal{I}^{N}_{i}$. By a slight abuse of notation, we also use the symbol $\mathcal{Q}^{N}$ for the operator $\mathcal{Q}^{N}:\mathbb{R}^{N\times N}\to L^{\infty}(\mathcal{I}\times\mathcal{I})$, mapping matrices $B\in\mathbb{R}^{N\times N}$ onto piecewise constant functions on $\mathcal{I}\times\mathcal{I}$,
@@ 3. Main assumptions
[…]
@@
Assumption (3.3) restricts the validity of our results to 010-1 valued graphons, i.e., $\mbox{range}(w)\subset\{0,1\}$. In this case one can show [19, Proposition 8.24] that the usual convergence of $w^{N}$ to $w$ in the cut norm indeed implies convergence in the $L^{1}$ norm topology. Of course, due to the uniform boundedness of $w^{N}$ in $L^{\infty}(\mathcal{I}\times\mathcal{I})$, the convergence (3.3) holds also in the norm topology of $L^{q}(\mathcal{I}\times\mathcal{I})$ for any $q<\infty$. Let us explicitly note that (3.3) excludes important classes of graphs. For instance, for the generic random Erdös-Rényi graph model [4], where pair of nodes are connected with probability $p\in(0,1)$, we have $w\equiv p$ on $\mathcal{I}\times\mathcal{I}$. Obviously, it is then not possible that $w^{N}\to w$ strongly in $L^{1}$, since the functions $w^{N}$ are piecewise constant taking only values 0 and $1$. We shall generalize our theory to include random graph models in a future work.
To be able to pass to the continuum limit in (1.3)-(1.4), we also need to assume that the edge lengths $\mathrm{L}^{N}=(\mathrm{L}^{N}_{ij})_{i,j=1}^{N}$ converge in an appropriate sense. In particular, for $\ell^{N}:=\mathcal{Q}^{N}[\mathrm{L}^{N}]$ we impose the assumption that
$\left\|\ell^{N}\right\|_{L^{\infty}(\mathcal{I}\times\mathcal{I})}=\sup_{i,j\in[N]}\mathrm{L}^{N}_{ij}\leq 1,$ (3.4)
@@
[…]
Assumption (3.4) is a natural assumption meaning that the sequence of graphs $\mathbb{G}^{N}$ is located within a compact set in the “physical space”. The particular bound $\mathrm{L}_{ij}\leq 1$ can be achieved through rescaling and is therefore assumed without loss of generality. A trivial way to ensure validity of assumption (3.5) is to fix $\ell\in L^{\infty}(\mathcal{I}\times\mathcal{I})$ and set $\ell^{N}:=\mathcal{Z}^{N}[\ell]$. The lengths $\mathrm{L}^{N}_{ij}$ are then the corresponding values of $\ell^{N}$, i.e., such that $\ell^{N}:=\mathcal{Q}^{N}[\mathrm{L}^{N}]$ holds. Again, due to the uniform boundedness of $\ell^{N}$ in $L^{\infty}(\mathcal{I}\times\mathcal{I})$, the limit in (3.5) holds also with respect to the norm topology of $L^{q}(\mathcal{I}\times\mathcal{I})$ for any $q<\infty$.
Assumptions (3.2)–(3.5) are sufficient for proving our first main result, which establishes the $\Gamma$-convergence of the sequence of (properly rescaled and reformulated) energy functionals. For our second main result, which states that the sequence of global minimizers of the discrete functionals (1.3)-(1.4) converges to the global minimizer of the limiting functional, we need to adopt one more assumption. Namely, we need to assume that $\gamma>2$ and that the sequence of reciprocal values $\left(\ell^{N}\right)^{-1}$ is uniformly bounded in $L^{\frac{2(\gamma+1)}{\gamma-2}}(\mathcal{I}\times\mathcal{I})$,
$\sup_{N\in\mathbb{N}}\left\|\left(\ell^{N}\right)^{-1}\right\|_{L^{\frac{2(\gamma+1)}{\gamma-2}}(\mathcal{I}\times\mathcal{I})}<\infty.$ (3.6)
Admittedly, this is a rather strong assumption and constitutes a key limitation of our approach. To gain an intuition about the satisfiability of (3.6), let us consider the case when the vertices in $\mathbb{V}^{N}$ are generated as random points in a convex compact set $K\subset\mathbb{R}^{d}$. I.e., we identify $\mathbb{V}^{N}$ with a cloud of $N\in\mathbb{N}$ points drawn independently from the uniform distribution on $K$. The edge lengths $\mathrm{L}_{ij}$ are then simply the Euclidean distances between the points (vertices) $i$ and $j$. The law of large numbers [2] gives then
$\lim_{N\to\infty}\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j\neq i}^{N}\mathrm{L}_{ij}^{-q}=\int_{0}^{\infty}s^{-q}f(s)\,\mathrm{d}s,$ (3.7)
where $f=f(s)$ is the probability density for the Euclidean distances of the point cloud. For the case when $K$ is a hypercube in $\mathbb{R}^{d}$, the probability density $f=f(s)$ is known explicitly; see, e.g., [16, 17] and references therein. For $d=2$ and $K=[0,1/\sqrt{2}]^{2}$, we have $\mathrm{L}_{ij}\leq 1$ and
@@
complemented by a more complicated formula for values of $s\in\left(\frac{1}{\sqrt{2}},1\right]$. However, $f(s)$ is decreasing for $s>1/\sqrt{2}$ (see [17]) so that we have the bound $f(s)<f\left(1/\sqrt{2}\right)=2\sqrt{2}(\pi-3)$ for $s>1/\sqrt{2}$. Of course, $f(s)=0$ for all $s>1$. Then, in (3.7) we have
@@
Obviously, the right-hand side is finite if and only if the first integral is finite, which holds if and only if $q<2$. Unfortunately, this makes assumption (3.6) unsatisfiable in dimension $d=2$, since
$\frac{2(\gamma+1)}{\gamma-2}>2\qquad\mbox{for all }\gamma>2.$
In general, the probability density for the $d$-dimensional cube behaves like $s^{d-1}$ as $s\to 0+$; see [17]. Consequently, (3.7) is finite if and only if $q<d$. Therefore, assumption (3.6) requires
$\frac{2(\gamma+1)}{\gamma-2}<d,$
which is equivalent to
$\gamma>\frac{2d+2}{d-2}=2+\frac{6}{d-2}.$
In particular, for $d=3$ we need $\gamma>8$, while for $d=4$ we need $\gamma>5$, and for $d=5$ we need $\gamma>4$.
@@ 4. Reformulation of the model and main results
For the purpose of deriving the graphon limit of the model (1.3)-(1.4), we first introduce a regularization by adding a fixed constant $r>0$ to all conductivities $C_{ij}$. This is necessary to ensure uniform ellipticity of the Kirchhoff law (1.3) and was also introduced in the previous works [9, 10] and [11] where continuum limits were derived. It is then convenient to introduce the new set of variables $B_{ij}:=\frac{C_{ij}+r}{\mathrm{L}_{ij}}$. Moreover, the energy functional $E^{N}$ has to be scaled by $1/N^{2}$, which can be interpreted as the reciprocal of the square of the “mean” edge length. This is due to the fact that we are embedding the graph $\mathbb{G}^{N}$, which is inherently a one-dimensional structure, into the two-dimensional interval $\mathcal{I}\times\mathcal{I}$. The same scaling was adopted in [9, 10, 11] in order to obtain a meaningful continuum limit. We therefore introduce the rescaled energy functional
$F^{N}[B]:=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\left(B_{ij}(P_{j}-P_{i})^{2}+\frac{\nu}{\gamma}B_{ij}^{\gamma}\left(\mathrm{L}_{ij}^{N}\right)^{\gamma+1}\right)W^{N}_{ij},$ (4.1)
@@
Clearly, we have $F^{N}[B]=E^{N}[C+r]/N^{2}$, with $(C+r)_{ij}:=C_{ij}+r$. We fix $r>0$ and, recalling the assumption $\mathrm{L}_{ij}\leq 1$, pose (4.1)–(4.2) on the convex set $\mathbb{B}^{N}_{r}$ defined in (2.4). We shall also use the “weak formulation” of (4.2), which is obtained by its multiplication by $\Phi_{i}$ for a “test vector” $\Phi\in\mathbb{R}_{0}^{N}$ and summation over $i\in[N]$,
@@
where we use the concise notation $\mathrm{d}w(x,y)$ for $w(x,y)\,\mathrm{d}x\,\mathrm{d}y$ and $w\in L^{\infty}(\mathcal{I}\times\mathcal{I})$ is given in (3.3). The functional $\mathcal{F}$ is defined on the set $L^{\infty}_{r}(\mathcal{I}\times\mathcal{I})$ introduced in (2.1). The scalar pressure $p=p[b]\in L^{2}_{0}(\mathcal{I})$ is obtained as a solution of the nonlocal diffusion (Poisson-type) equation
@@
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)(\overline{\nabla}p(x,y))(\overline{\nabla}\varphi(x,y))\,\mathrm{d}w(x,y)=\int_{\mathcal{I}}\sigma(x)\varphi(x)\,\mathrm{d}x\quad\forall\varphi\in L^{2}_{0}(\mathcal{I}),$ (4.5)
where we used the notation (2.5). Let us note that the functional (4.4)–(4.5) is invariant with respect to any measure preserving change of variable on the interval $\mathcal{I}$. Indeed, if $\psi:\mathcal{I}\to\mathcal{I}$ is any such measure preserving bijection, then the value of the energy remains the same for the transformed functions $\tilde{b}(x,y):=b(\psi(x),\psi(y))$, $\tilde{\ell}(x,y):=\ell(\psi(x),\psi(y))$, $\tilde{w}(x,y):=w(\psi(x),\psi(y))$ and $\tilde{p}(x):=p(\psi(x))$, $\tilde{\sigma}(x):=\sigma(\psi(x))$. This invariance is important since $w$ and $\tilde{w}$ represent the same graphon [3, 19]. An analogous invariance obviously also holds for the discrete model (4.1)–(4.2) with respect to permutations of the set of indices $[N]$.
@@
$\mathcal{F}^{N}[b]:=\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)\left(\overline{\nabla}p^{N}(x,y)\right)^{2}+\frac{\nu}{\gamma}b(x,y)^{\gamma}\ell^{N}(x,y)^{\gamma+1}\,\mathrm{d}w^{N}(x,y),$ (4.6)
with $\ell^{N}=\mathcal{Q}^{N}[\mathrm{L}^{N}]$, $w^{N}=\mathcal{Q}^{N}[W^{N}]$, and $p^{N}\in L^{2}_{0,N}(\mathcal{I})$ is the unique piecewise constant solution of the approximate Poisson equation
$\frac{1}{2}\int_{\mathcal{I}}\int_{\mathcal{I}}b(x,y)(\overline{\nabla}p^{N}(x,y))(\overline{\nabla}\varphi^{N}(x,y))\,\mathrm{d}w^{N}(x,y)=\int_{\mathcal{I}}\sigma(x)\varphi^{N}(x)\,\mathrm{d}x\quad\forall\varphi^{N}\in L^{2}_{0,N}(\mathcal{I}).$ (4.7)
@@
Moreover, the sequence of functionals $(\mathcal{F}^{N})_{N\in\mathbb{N}}$, given by (4.6)–(4.7), $\Gamma$-converges in $L_{r}^{\omega}(\mathcal{I}\times\mathcal{I})$, $\omega=\max\{2,\gamma\}$, to the functional $\mathcal{F}$ given by (4.4)–(4.5), in the following sense: